find the equation of a line with y intercept 2 and point 4,4
Answer by nerdybill(7384)

find the equation of a line with y intercept 2 and point 4,4
.
Plug in the given point (4,4) and the y-intercept into:
y = mx + b
and solve for m (the slope).
.
4 = 4m + 2
2 = 4m
1/2 = m
.
At this point, we now have:
y intercept at 2
and
m = 1/2
.
So, plug it all back into:
y = mx + b
y = (1/2)x + 2
or
y = x/2 + 2